A full-service restaurant is considering opening a new facility in a specific city. The table below shows its ratings of four factors at each of two potential sites. 

Factor
WeightGary Mall.Belt Line
Affluence of local population203030
Traffic flow.402020
Parking availability203020
Growth potential.201030

 The score for Gary Mall is_______  and the score for Belt Line is _______ 

 A. 34:28 

 B. 22:24 

 C. 18; 120 

D. 90; 100 

 E. None of the above.
Answer #1

B. 22, 24

Weighted score for Gary Mall = (0.2*30 + 0.4*20 + 0.2*30+ 0.2*10)

= (6+8+6+2)

= 22

Weighted score for Belt Line = (0.2*30 + 0.4*20 + 0.2*20 + 0.2*30)

= (6+8+4+6)

= 24